The near point (the smallest distance at which an object can beseen clearly) and the far point (the largest distance at which anobject can be seen clearly) are measured for six differentpeople.

near point
({\rm cm})
far point
({\rm cm})
Avishka
40
\infty
Berenice
30
300
Chadwick
25
500
Danya
25
\infty
Edouard
80
200
Francesca
50
\infty
a) Which, if any, of these people are nearsighted(myopic)?
b) Which, if any, of these people require bifocals to correcttheir vision?
c) Which, if any, of these people's vision can be correctedusing only converging lenses?
Answer #1
Concepts and reason

The concept required to solve the given problem is far point and near point.

Use the definition of near point and far point to conclude who amongst the listed people are nearsighted, require bifocals and converging lens.

Fundamentals

Near Point is the smallest distance at which an object can be clearly seen by a normal eye. It is 25cm25{\rm{ cm}} for a normal eye.

Far Point is the largest distance at which an object can be seen clearly by a normal eye. It is \infty for a normal eye.

(a)

The near point that is the smallest distance at which an object can be clearly seen by a normal eye is 25cm25{\rm{ cm}}.

The far point that is the largest distance at which an object can be seen clearly by a normal eye is \infty .

From the table given Berenice, Chadwick and Edouard has far-point less than infinity due to which they are nearsighted.

(b)

Bernice and Edouard have near point greater than 25cm25{\rm{ cm}} and far point less than \infty due to which they suffer both from myopia and hyperopia.

Due to the reason mentioned above the listed people will require bifocals to correct vision.

(c)

The near point that is the smallest distance at which an object can be clearly seen by a normal eye is25cm25{\rm{ cm}}.

The far point that is the largest distance at which an object can be seen clearly by a normal eye is\infty .

A hyperopic eye can see far off things clearly but the near point for a hyperopic eye increases due to which it is greater than 25cm25{\rm{ cm}} for that eye.

From the table given Avishka and Francesca has farsightedness due to which their near point is greater than25cm25{\rm{ cm}}.

Ans: Part a

Berenice, Chadwick and Edouard are nearsighted.